What is the expected freezing point of a 0.50 m solution of Na2CO3 in water? If for water is 1.86 c/m
Answer #1

Depression in freezing point, ∆T = iKm

For Na2CO3, i = 3 , m = 0.50 , K = 1.86

∆T = 3×1.86×0.50 = 2.79°C

Freezing point = 0 - 2.79 °C = -2.79°C

( 0°C is the freezing point of water).

Please give feedback!